Tech Staff Meeting Summary 3/6/02

Current Top Priorities are:ESEE, Complete DBB, The Big Move

CVTN is running. Need one more visit to get comm going.

CBHT comm. Switch combined with dramatic k2ew config changes appeared to work. Steve got said a new firmware upgrade may also solve the problem.

HDBT. Need a fence and gravel for ground cover.

NAIT. NWS does not know when the VPN will be installed. Mitch will ask the NWS Houston contact for more info.

CUET. Provided additional information to Pat Bohannan at TEMA regarding the software connections. He's working on the firewall issues (we hope).

SFTN. K2 and MSS100 are fine. No comm out the ethernet. Needs debugging.

TUMT. Mitch reconfigured all k2ew modules to be similar to CBHT (i.e. more robust with dropouts). Seems to have fixed TUMT.

DBB. Steve suggested a terminal server/watchdog to do dynamic resets when the hang. Has been working great though so this is backburner (need to order a bigger backburner and more pilot lights). ISTI worm module to monitor outages should best wait until RPC is running. Jim will check into getting the dish hung at NMAD and LNXT.

Jim got some additional quotes for the NMAD fence. It is in Purchasing.

GPS. Still no detail in GPS status. Still can't fix what we don't know is broke. Mitch insists on more than just a binary list of station outage. GPS personnel need to investigate as much as possible from here prior to making mainenance runs.

Greg finished the box for the sensors in the yard. The MSS4 is in the computer room. Scream will not talk directly to the lantronix box (it wants either a serial port or another scream) Mitch installed the lantronix software on darkside to make the MSS4 ports appear as virtual serial but had trouble. Need to verify comm with the MSS4 out of the loop, then revisit the virtual port issue.

ET. John has been spending alot of time on PII boxes, destroyed and moved stations, etc. Need to discuss when John gets here. Need to closely look at what is really needed in NC. Mitch is rerouting the internet communications to try to reduce the load. Need to make a trip over there to help get things running. Need a couple aironets at ATHN to get it online but not until we know whether or not they will be moving (should know soon). Comm at ATHN is down. John is removing the bsd boxes and sending the freewaves to Memphis. Steve will reconfigure and setup to get NT to NT ppp going directly from athn EMA to the star mountain earthworm. SWET is also down; need to replace the sensor.

NM. With one notable exception all New Madrid stations are up (GNAR had a crack in the conduit connection between the inner and outer borehole casings due to differential heaving; ain't the embayment muck wonderful). HAWT was repaired and is down again already (likely shotgun).

Chapman and Munsey are working on a site for Oakridge to put a CMG3 or 40T. Mitch will provide more info on proposed KY NSN station to get contacts.

FDL. Russ has the one broken K2's and 3 others need config. Using aftershock params from Paul as "standard" config to say they are ready to go. Any customization will need to be done just prior to deployment. Also need to do "end to end" tests.

NHIN. Filson made a personal phone call to thank us for the NHIN installation. Steve cleaned up the display alot. Mitch is working on better index files for the helicorders. Surface sensors are very culturally noisy but borehole shows promise.

PII. Most of PI has been cleaned out. Still need to go through PII but it is in relatively good shape. Need to get going on tuning and calibrating.

Monroe. Due to telemetry concerns of other sites, need to design for original capable of keeping its "head above water."

RPC. Mitch has acquired funds for the Regional Processing Center hardware (RPC). Greg, Steve, and Mitch are working on power specs etc to get the racks built.

FY02 ANSS. Will be bringing techs from USC, SLU, VPI, and Parker to Memphis. Don't know what we're installing so we must assume it won't be too much different from last year. Meeting will be March 25 and 26 but unfortuanely Cannon probably can't make it.

Greg is working on version 101f of the I-40 installation.

SFTN needs solar backup power. Done.

Need to do site visits of Lamont stations sometime this year.

The big move. To better accomadate administrative archives we need to do the following. Rydelek to 3904/207, Greganti to Rydelek, Spikes to Greganti, Active records in Archive Room to Greganti, Archive records in attic and elsewhere to Archive room. Need to be careful how we do Rydelek since he's out of town. Will need to do prep work then get contract assistance to do a monster one day move.

3904/215E. Will expand into the closet and be used as a computer setup/testing lab.

Need to get cost estimates for a vehicle and generator.

Propane tank at LNXT has been retrofitted.

Need to organize a Giddings workshop this summer.

Kaiser mud is gone. Paul gets the bags.

Need to replace Parker's Suburban.

Need to get moving on siting for ESEE and ANSS. Mitch will get ESEE lat/lon targets and provide ANSS targets. Need to interface with Chuck on the ESEE sites. Pat has begun repair on S-13's.

Need to assist Chiu/UKY in profiling experiment. Just a day or two to show them where things are. UKY is responsible for doing the actual experiment.